OpenStudy (anonymous):

kumar can paint his room in 10hrs. delia can paint the same room in 8hrs. to the nearest minute, how long will it take if they paint the room together????? answers f. 4h 4min. g. 4h 27min h. 6h 22min j. 9h

OpenStudy (anonymous):

\[\frac{8\times 10}{8+10}\] you can do the conversion

OpenStudy (anonymous):

so i add them and divide and then convert?

OpenStudy (anonymous):

yes it is \[\frac{80}{18}\] which reduces easily

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

i think you get \(4\tfrac{4}{9}\) which you then apparently have to convert to minuutes maybe for hours and 28 minutes? 27 minutes? something like that